OpenStudy (anonymous):

Find an explicit rule for the nth term of a geometric sequence where the second and fifth terms are -12 and 768, respectively.

OpenStudy (anonymous):

i understand how to do this when the two terms are already consecutive, but im not sure where to start with this one

OpenStudy (jhannybean):

what is the ratio?

OpenStudy (anonymous):

how do i find the ratio without two consecutive terms?

OpenStudy (jhannybean):

Grr...

OpenStudy (anonymous):

:x sorrry

OpenStudy (anonymous):

im not just looking for an answer, i legitimately want help understanding how to work the problem, i just need help through most of it

OpenStudy (jhannybean):

Yeah..i'm trying to work it out to explain it to you,but having difficulty finding the ratio.

OpenStudy (radar):

Would knowing the ratio r be helpful?

jimthompson5910 (jim_thompson5910):

a2 = -12 a3 = a2*r = -12r a4 = a3*r = (-12r)*r = -12r^2 a5 = a4*r = (-12r^2)*r = -12r^3

jimthompson5910 (jim_thompson5910):

because a5 = 768, we can say 768 = -12r^3

OpenStudy (radar):

There you go, now you can solve for r

OpenStudy (jhannybean):

nice!

OpenStudy (anonymous):

so, 768/-12 = r^3 -64 = r^3 -4 = r

OpenStudy (anonymous):

an = a1(-4)^n-1

OpenStudy (anonymous):

im sorry, but now now how would i find the first term to use in the rule?

OpenStudy (loser66):

show me how you have the form above? an? one more question, does your prof ask you have to prove what you claim before using it to find our the nth term?

OpenStudy (loser66):

*out

OpenStudy (anonymous):

a subscript n

jimthompson5910 (jim_thompson5910):

Use the idea Second term = (First term)*r

OpenStudy (anonymous):

nth term = (first term)(ratio)^n-1

OpenStudy (anonymous):

that was the basis of my lesson. im virtual schooled, so the lessons are very impersonal, leaving much to the student to figure out on their own

OpenStudy (jhannybean):

\[\large a_{2} =?\]\[\large r=?\]\[\large a_{1}=?\] plug into equation \[\large a_{2}=a_{1}*r\]

OpenStudy (loser66):

or a1 = a2/r

OpenStudy (anonymous):

would it be an = 3(-4)^n-1

OpenStudy (jhannybean):

yep

OpenStudy (anonymous):

okay, thank you all